OpenStudy (anonymous):

help please !!! y=x^2+5x+6 i need to know te vertex and ordered pair???

OpenStudy (anonymous):

the x coordinate of the vertex of the parabola: \(\large y=ax^2+bx+c \) is \(\large h=-\frac{b}{2a} \) to find the y coordinate of the vertex, find f(h)

OpenStudy (anonymous):

is the vertex -2.5,4.75

OpenStudy (anonymous):

dpalnc is the vertex -2.5, 4.75?

OpenStudy (anonymous):

your x coordinate is correct, y coordinate is wrong.

OpenStudy (anonymous):

so what would be the y coordinate? and the ordered pair? please help

OpenStudy (anonymous):

can anyone please help me I need to know the ordered pair and the vertex